MATRIX ISOLATION INFRARED OBSERVATION OF THE HSiN_{2} AND H_{2}SiN_{2} COMPLEXES USING A 121 NM VACUUM-ULTRAVIOLET PHOTOLYSIS SOURCE

\underline{JAY C. AMICANGELO}, CHRISTOPHER T. DINE, DANIEL G. IRWIN, CYNTHIA J. LEE, NANCY L. SAXTON, School of Science, Penn State Erie, Erie, PA 16563.

Matrix isolation infrared spectroscopy was used to characterize the complexes of H_{2}Si and HSi with N_{2} that result from the vacuum-ultraviolet photolysis of silane (SiH_{4}) in cryogenic nitrogen (N_{2}) matrices. Experiments were performed by depositing mixtures of SiH_{4} with N_{2} onto a CsI window at 12 K while simultaneously photolyzing the mixture with 121 nm vacuum-ultraviolet radiation from a hydrogen resonance lamp. The infrared bands of the H_{2}SiN_{2} complex observed in these experiments are the N-N stretching mode at 2274.2 cm^{−1} and the H_{2}Si asymmetric and symmetric stretching modes at 2013.6 and 2009.4 cm^{−1}, respectively. The infrared bands of the HSiN_{2} complex observed in these experiments are the N-N stretching mode at 2023.9 cm^{−1}, the Si-H stretching mode at 2006.6 cm^{−1}, and the H-Si-N bending mode at 813.7 cm^{−1}. The assignment of these bands to the H_{2}SiN_{2} and HSiN_{2} complexes is established by performing experiments with isotopic reagents (SiD_{4}, ^{15}N_{2}), by performing matrix annealing experiments (warming to 20 - 30 K and refreezing to 12 K), by performing mercury-xenon lamp matrix photolysis experiments, and by comparison to density functional theory calculations at the B3LYP/aug-cc-pVTZ level.